OpenStudy (anonymous):

Given the functions k(x) = 5x − 8 and p(x) = x − 4, solve k[p(x)] and select the correct answer below. k[p(x)] = 5x − 12 k[p(x)] = 5x − 28 k[p(x)] = 5x2 − 12 k[p(x)] = 5x2 − 28

OpenStudy (haseeb96):

if f(x)= 2x^2 and g(x)=3x now f[g(x)] = f[3x] = 2(3x)^2 =18 x^2 is your answer now solve your question same like this one

OpenStudy (anonymous):

so would the answer be D?

OpenStudy (lynfran):

no

OpenStudy (lynfran):

k[p(x)]=5(x-4)-8

OpenStudy (lynfran):

multiply out the bracket and simplify ...

OpenStudy (lynfran):

what do we get @kkailyn

OpenStudy (haseeb96):

no B is correct answer

OpenStudy (anonymous):

B?

OpenStudy (lynfran):

yes B is correct

OpenStudy (anonymous):

@LynFran , thank you so much, the formula you showed just taught me how to solve this type of problem correctly